OpenStudy (anonymous):

If f(x) = 2x + 1 and g(x) = −2x − 3, find (f + g)(−1).

OpenStudy (anonymous):

That just means, add f(x) + g(x) together. then evaluate that at x = -1

OpenStudy (anonymous):

is - 2

OpenStudy (anonymous):

yep. correct.

OpenStudy (anonymous):

can u show the work please kinda confused

OpenStudy (anonymous):

(f+g)(x) = f(x) +g(x) = (2x+1)+(-2x-3) = 0x -2 = 2 does that make sense

OpenStudy (anonymous):

-2 sorry

OpenStudy (anonymous):

yes thnak you so its not negative 2 its postive 2

OpenStudy (anonymous):

yes its negative two. Thats what i meant